# Boosting Open-Domain Continual Learning via Leveraging Intra-domain Category-aware Prototype

### v1

@@ Abstract
The task of open-domain continual learning in Vision-Language Model (VLM) has become both a practical and yet challenging problem. In the Open-Domain Continual Learning (ODCL) task, VLMs are required to incrementally learn both seen and unseen categories from diverse datasets, enhancing the recognition of new categories while maintaining the ability to recognize previously seen categories. As Contrastive LanguageImage Pre-training (CLIP) is a VLM trained on a large number of image-text pairs, it possesses strong zero-shot extraction capabilities for features across different domains. However, existing methods addressing the ODCL task often overlook the inherent domain gaps. In this work, we extract domain prior information using a training-free method to enhance the performance of CLIP, named DPeCLIP. Specifically, we leverage the strong domain information extraction capabilities of the CLIP model to achieve training-free Domain Prior Classifier across different domains. Additionally, we introduce the Domain Prior Prompt method, which incorporates domain prior information into the model to bridge the significant gaps between domains. Extensive experiments conducted on 11 datasets demonstrate that our approach outperforms the current state-of-the-art methods, validating its effectiveness.
Figure 1: (a) Comparison of two approaches for solving the ODCL-CIL task: one approach involves using all seen categories for the classification of test images, while the other method employs a Task-ID discriminator to determine the domain corresponding to the test image, and then selects the categories of the corresponding domain for classification. (b) Comparison of Task-ID classification accuracy between MoE-Adapters and our method. (c) Comparison of the performance of our method and other methods in ODCL-CIL task on the Caltech101 and CIFAR100 datasets. The ODCL task requires evaluating both seen and unseen datasets. The black dashed line on the left indicates that the test dataset has not been trained on, and it is assessed through the model’s zero-shot capability.
@@ Introduction
Continual learning (CL) (dhar2019learning; rebuffi2017icarl) aims to enable models to continually acquire new knowledge, such as learning to recognize new object categories, while addressing the catastrophic forgetting of old knowledge. In recent years, CL has gained significant attention and has been applied in various fields, including autonomous driving and medical diagnosis. However, traditional continual learning methods are largely limited to single-domain data, failing to effectively address the complexities of real-world scenarios. In practice, data often comes from multiple domains, and there may be significant differences between these domains.
@@
To address this issue, open-domain continual learning (ODCL)  (zheng2023preventing; li2024coleclip) has been explored by many recent researchers (zheng2023preventing; yu2024boosting; li2024coleclip). The ODCL task utilizes Vision-Language Model (VLM) (jia2021scaling; li2022blip; radford2021learning) to linearly learn data from different domains, acquiring new domain knowledge while preventing the forgetting of old domain knowledge. Due to the zero-shot classification capabilities of VLM, ODCL is tested on both seen and unseen domains, necessitating the preservation of VLM’s original zero-shot classification abilities. Similar to traditional continual learning, the ODCL task includes both class-incremental (CIL) and task-incremental (TIL) settings. Compared to ODCL-TIL settings, ODCL-CIL does not have access to the Task-ID of the current test image during inference.
@@
ODCL task faces two significant challenges: (1) Two types of Forgetting: Unlike traditional CL task, ODCL task faces not only the forgetting of old knowledge but also the forgetting of zeroshot capabilities. (2) Influence of Category-relatedness Between Domains: For example, although ”wheelchair” in Caltech101 and ”chair” in CIFAR100 belong to entirely different categories, their semantic similarity can cause confusion in the ODCL-CIL task, leading to a decrease in dataset performance.
@@
Currently, there are two primary technical approaches to addressing ODCL task, as illustrated in Figure 1(a). For the first approach, during the inference stage, the model utilizes the embeddings of all seen categories to classify the test image regardless of which domain the test image comes from. ZSCL (zheng2023preventing) leverages a large reference dataset combined with the distillation of original CLIP knowledge. Despite the distillation, fully fine-tuning CLIP results in significant forgetting. ColeCLIP (li2024coleclip) tries to mitigate the interference of similar categories across different domains by maintaining a large vocabulary and employing a finely designed task prompt. But, since ColeCLIP utilizes the same text embedding for the same category across different datasets (domains), although momentum is used to slow down changes in text embeddings during this process, the lack of updates to the old task prompt leads to inconsistency between the text embedding and the old task prompt. As shown in Figure 1(c), with the progression of training stages, ColeCLIP exhibits significant catastrophic forgetting issues. Besides, its zeroshot performance also declines. These methods try to alleviate the forgetting problem, but since the model cannot access data from multiple domains simultaneously during training, and similar categories from different domains can cause confusion during testing, addressing the category-relatedness problem becomes very challenging.
@@
The second approach involves first determining the Task-ID of the test image using a Task-ID discriminator, and then classifying it using the category set corresponding to the identified Task-ID. For example, MoE-Adapters (yu2024boosting) first propose a combination of Auto-Encoder (AE) and AlexNet (krizhevsky2012imagenet) to determine the Task-ID followed by exploring the performance of CLIP on ODCL task by adopting a Mixture of Experts (MoE) (jacobs1991adaptive) structure. Despite the expert freezing mechanism, the continual updating of experts still results in knowledge forgetting. As shown in Figure 1(b) and (c), the effectiveness of the Task-ID discriminator directly affects the model’s final performance. Due to the limited performance of the Task-ID discriminator, the overall performance of the model is affected.
@@
Compared to the first approach, the second approach offers a better solution to addressing the issue of the category-relatedness, namely that a more robust Task-ID discriminator can more effectively resolve this problem. Besides, we find that even with a finely designed domain prompt, updating it using cross-domain information still fails to alleviate the forgetting issue (li2024coleclip), since the task prompt can still be influenced by the inconsistency between text embeddings and old task prompt. Therefore, if the following two factors can be considered, it will greatly solve the category-relatedness problem and mitigate the forgetting of old domains: (1) correctly identifying the Task-ID of a test image, and using only the category set corresponding to the Task-ID, and (2) preserving the knowledge only related to each domain. Additionally, the model should retain the original knowledge of CLIP to maintain a good zeroshot capability.
@@
Based on the above considerations, we propose a simple yet effective solution: leveraging intra-domain category-aware prototypes for ODCL in CLIP (radford2021learning) framework, dubbed as Domain Prototype enhanced CLIP (DPeCLIP). To maintain the original zeroshot capability of CLIP, we propose to average the original outputs of CLIP’s image and text modalities belonging to the same category within a domain as prototypes. We propose to (1) distinguish the Task-ID of a test image, and (2) preserve the knowledge only related to each domain, where prototypes is the key to bridging them. Firstly, we propose a robust, training-free Task-ID discriminator, by utilizing prototypes as classifiers for identifying Task-IDs. Compared to MoE-Adapters, our method significantly improves Task-ID accuracy without introducing additional training parameters. Secondly, to maintain the knowledge corresponding to each domain during training, we incorporate intra-domain category-aware prototypes as domain prior prompts into the training process. For the text branch, we propose a text self-attention module which encodes the relationships of categories within a domain as prompts into the text branch. For the image branch, unlike the text branch, we introduce an image cross-attention module which uses instance-level image embeddings to query category-level prototypes. This ensures that the instance prompt carries information about the relationships between various categories, enabling the learned prompt to encapsulate both domain and instance information.
@@
  • We propose a robust, training-free Task-ID discriminator method that utilizes domain-specific, category-aware prototypes as classifiers to effectively distinguish test images from the original domain. This method not only addresses the issue of strong category-relatedness across different domains, but also significantly improves the model’s performance on ODCL task.
  • We incorporate category-domain class-aware prototypes as domain prior prompts into the training process to maintain the knowledge corresponding to each domain. This method not only enhances the model’s recognition capability for new categories, but also maintains good zero-shot classification ability.
  • Through comprehensive experiments on 11 datasets, we demonstrate that our method achieves state-of-the-art (SOTA) performance in both ODCL-CIL and ODCL-TIL settings, compared to previous best methods.
@@
To address the issue of catastrophic forgetting, continual learning can be broadly categorized into three primary approaches: replay-based methods (bang2021rainbow; chaudhry2018riemannian; rebuffi2017icarl; lopez2017gradient; prabhu2020gdumb; shin2017continual), regularization-based methods (kirkpatrick2017overcoming; zenke2017continual; li2017learning; dhar2019learning; douillard2020podnet; hou2019learning), and parameter expansion-based methods (gao2023unified; smith2023coda; wang2022foster; yan2021dynamically; zhou2022model; wang2022dualprompt; wang2022learning; wang2022s). Replay-based methods maintain a buffer of previously encountered data and leverage this replay buffer during the learning of new data to retain the original knowledge. Regularization-based methods aim to preserve prior knowledge by constraining the direction of model updates, with knowledge distillation (hinton2015distilling) being a prevalent technique. Parameter expansion-based methods mitigate forgetting by expanding the model’s architecture to accommodate new tasks. Recently, several parameter-efficient fine-tuning (PEFT) based on parameter expansion methods have emerged, such as Low-Rank Adaptation (LoRA) (hu2021lora), prompt-based methods (lester2021power), and adapter modules (houlsby2019parameter). L2P (wang2022learning) addresses the issue of preserving old knowledge by maintaining a large prompt pool. DualPrompt (wang2022dualprompt) further mitigates catastrophic forgetting by employing both general and specific prompts. CODA-Prompt (smith2023coda) extends this approach by increasing the scale of prompts and utilizing more prompts simultaneously through a weighting mechanism. LAE (gao2023unified) explores the use of three PEFT methods by implementing both online and offline update strategies.
@@
Currently, numerous studies focus on the performance of CLIP in continual learning, which can be divided into two categories: traditional continual learning (ding2022don; zhou2023learning; wang2022s) and ODCL tasks (zheng2023preventing; yu2024boosting; li2024coleclip).
For traditional continual learning with CLIP, VR-LwF (ding2022don) retains old knowledge through a simple vocabulary distillation method, Proof (zhou2023learning) employs a cross-modal fusion module to fully utilize knowledge and task-specific mapping layers to alleviate forgetting, and S-Prompt (wang2022s) uses k-means to save the center of each domain as a Task-ID discriminator, selecting the corresponding domain’s prompt for test images.
ZSCL (zheng2023preventing), as the first method proposed for open-domain tasks, balances newly learned knowledge and old knowledge through feature-level distillation with an external dataset. MoE-Adapters (yu2024boosting) improve performance in ODCL task by introducing an external dataset to train corresponding autoencoders for Task-ID identification and using a MoE structure. CoLeCLIP (li2024coleclip) extends open-domain tasks, by maintaining a vocabulary and designing skillfully task prompts to store task-specific knowledge, but this approach inevitably leads to performance degradation in ODCL-CIL task. In comparison, our method achieves significant performance improvement by determining the Task-ID through a Domian Prior Classifier approach.
Figure 2: Domain Prototype enhanced CLIP (DPeCLIP) framework.The three stages are presented as follows: (a) Prototype Calculation Stage, where category-aware prototypes are extracted using the original CLIP model. (b) Training Stage, where we utilize two sub-modules, Text Self-Attention (TSA) and Image Cross-Attention (ICA), to provide Domain Prior Prompts for different categories and images, conditioned on the domain prototypes as input. Additionally, we extract domain-level shared information using general prompts. (c) Inference Stage, where we use the Domain Prototypes to determine the Task-ID for test images and employ the corresponding Domain Components and category sets associated with the Task-ID for classifying the test images.
@@
CLIP is a model endowed with zero-shot classification capabilities, comprising an image encoder $\mathcal{I}(\cdot)$ and a text encoder $\mathcal{T}(\cdot)$. The image encoder can be configured as either ViT (dosovitskiy2020image) or ResNet (he2016deep), whereas the text encoder is based on the Transformer architecture (vaswani2017attention). These encoders are trained using contrastive loss on a vast amount of paired image-text samples, facilitating a high degree of alignment between image and text features, which imparts robust zero-shot classification abilities to the model.
To employ the CLIP model for classification tasks, an image $x$ is first processed through the image encoder to extract image features. The target classes $\{c^{1},c^{2},\cdots,c^{J}\}$ for discrimination are converted into corresponding textual descriptions $\{\hat{c}^{1},\hat{c}^{2},\cdots,\hat{c}^{J}\}$ using a template such as “a photo of a”, where $J$ represents the total number of categories. These textual descriptions are then fed into CLIP text encoder to generate text features. The cosine similarity between the text features and the image features is computed, and the candidate category with the highest similarity is selected as the final classification result.
@@
For the ODCL tasks, the model is required to sequentially learn from $N$ distinct datasets (domains). The primary objective of this task is to enhance the recognition of new categories while maintaining the ability to recognize previously learned categories. Given $N$ datasets, $\{\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{N}\}$, as the model learns these datasets linearly, after learning the $n$th dataset $\mathcal{D}_{n}$, it exhibits good classification capability on the previously learned datasets $\{\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{n}\}$. At the same time, the model aims to preserve its zero-shot classification ability for the unlearned datasets $\{\mathcal{D}_{n+1},\mathcal{D}_{n+2},\cdots,\mathcal{D}_{N}\}$.
To thoroughly evaluate the model’s performance on ODCL, there are two settings for inference: task-incremental learning and class-incremental learning, denoted by ODCL-TIL and ODCL-CIL, respectively. In the ODCL-TIL task, the Task-ID is known during inference, allowing us to utilize the category set corresponding to the test image. In contrast, in the ODCL-CIL task, the Task-ID is not accessible during inference, necessitating the use of all seen categories to classify the test image. Clearly, for ODCL-CIL task, the inability to obtain the task ID makes it a more challenging task compared to ODCL-TIL.
@@
Figure 2 illustrates the overall architecture of our method. Our method consists of three parts: (a) Prototypes Calculation Stage, (b) Training Stage, and (c) Inference Stage. In the Prototypes Calculation Stage, we utilize the CLIP image encoder to extract the image features of training images belonging to $j$th category within $n$th domain and average them to obtain the mean image feature $\bar{I}_{n}^{j}$. Likewise, we use the CLIP text encoder to extract the text features, combined with CLIP templates, such as “a photo of a”. By averaging these text features, we obtain the average text feature $\bar{T}_{n}^{j}$. Finally, by performing element-wise summation of $\bar{I}_{n}^{j}$ and $\bar{T}_{n}^{j}$, we derive one intra-domain category-aware Prototype $P_{n}^{j}$.
During the training stage, we propose Domain Prior Prompts through two sub-modules: Text Self-Attention (TSA) and Image Cross-Attention (ICA), conditioned on the prototypes as input. In addition, we also learn the shared information for each domain through the General Prompts. In this process, both domain prior prompt and the general prompt are updated using cross-entropy loss.
During the inference stage, we first utilize the domain prototypes to determine the Task-ID. Then, based on the identified Task-ID, we select the corresponding TSA, ICA modules, General Prompts, and the category set associated with the Task-ID from the domain prior prompt pool, which contains all trained components, to classify the test image.
@@
To obtain intra-domain category-aware prototypes for ODCL, we extract prototypes from the training set using the original CLIP model, as shown in Figure 2(a). Specifically, for the $j$th category in $\mathcal{D}_{n}$, containing $M$ training images, the image prototype $\bar{I}_{n}^{j}$ can be calculated via:
$\bar{I}_{n}^{j}=\frac{\sum_{m=1}^{M}\mathcal{I}(x_{n}^{j,m})}{M},$ (1)
where $x_{n}^{j,m}$ means the $m$th image belonging to the $j$th category of $\mathcal{D}_{n}$. We adopt different templates provided by CLIP to generate various textual descriptions for a category name. For the $j$th category in the $n$th domain $c_{n}^{j}$, the textual description generated by the $z$th template is denoted as $\hat{c}_{n}^{j,z}$. Then the text prototype $\bar{T}_{n}^{j}$ is calculated by:
$\bar{T}_{n}^{j}=\frac{\sum_{z=1}^{Z}\mathcal{T}(\hat{c}_{n}^{j,z})}{Z},$ (2)
where $Z$ represents the total number of templates.
Subsequently, we obtain the intra-domain category-aware prototypes $P_{n}^{j}$ by performing element-wise summation of $\bar{I}_{n}^{j}$ and $\bar{T}_{n}^{j}$.
@@
Moreover, we propose a category-aware domain prior prompt for the text encoder. Specifically, conditioned on the prototype $P_{n}=[{P_{n}^{1}}^{\top},{P_{n}^{2}}^{\top},\cdots,{P_{n}^{S}}^{\top}]^{\top}$ as input, where $P_{n}\in{\mathbb{R}^{S\times d^{\mathcal{T}}}}$. We propose a Text Self-Attention (TSA) module to generate the domain prior prompt $v_{p}$, $v_{p}\in{\mathbb{R}^{S\times d^{\mathcal{T}}}}$. The TSA module is depicted in Figure 2 (with Add and Norm omitted). We use the domain prototype $P_{n}$ as the input for ${Q}$, ${K}$, and ${V}$ in TSA, thereby obtaining $v_{p}$.
$Q=W_{Q}\cdot P_{n},K=W_{K}\cdot P_{n},V=W_{V}\cdot P_{n},$ (4)
@@
These domain prior prompts $v_{p}$ then replace the general prompt in the $h$-th layer of the text encoder. Thus, the output of the $h$th text encoder layer is:
@@
As shown in Figure 2(b), similar to the text branch, we use the general prompt to learn the general information of the domain for image encoder $\mathcal{I}(\cdot)$. In addition, we utilize the ICA module to learn the instance-level domain prior prompt, thereby providing both domain and instance information simultaneously. Specifically, the image passes through the patch embedding layer to obtain image tokens $[cls^{(1)},i_{1}^{(1)},\cdots,i_{l^{\prime}}^{(1)}]$, where $cls^{(1)}$ is class token in the first layer and subscript $l^{\prime}$ is the number of image tokens. In the first $h-1$ layers of the image encoder, we concatenate a general prompt $e_{g}$, where $e_{g}\in{\mathbb{R}^{1\times d^{\mathcal{I}}}}$ with the image tokens at each layer, and $d^{\mathcal{I}}$ means the dimension of image token. For example, for the $h-1$ layer $\mathcal{I}^{(h-1)}(\cdot)$ of image encoder, the general prompt $e_{g}^{(h-1)}$ is used by following formula.
$\begin{split}&[cls^{(h)},t_{1}^{(h)},\cdots,t_{l^{\prime}}^{(h)},e_{g}^{(h)}]=\\ &\mathcal{I}^{(h-1)}([cls^{(h-1)},t_{2}^{(h-1)},\cdots,t_{l^{\prime}}^{(h-1)},e_{g}^{(h-1)}])\end{split}$ (8)

### v2

@@ Abstract
Despite recent progress in enhancing the efficacy of Open-Domain Continual Learning (ODCL) in Vision-Language Models (VLM), failing to (1) correctly identify the Task-ID of a test image and (2) use only the category set corresponding to the Task-ID, while preserving the knowledge related to each domain, cannot address the two primary challenges of ODCL: forgetting old knowledge and maintaining zero-shot capabilities, as well as the confusions caused by category-relatedness between domains. In this paper, we propose a simple yet effective solution: leveraging intra-domain category-aware prototypes for ODCL in CLIP (DPeCLIP), where the prototype is the key to bridging the above two processes. Concretely, we propose a training-free Task-ID discriminator method, by utilizing prototypes as classifiers for identifying Task-IDs. Furthermore, to maintain the knowledge corresponding to each domain, we incorporate intra-domain category-aware prototypes as domain prior prompts into the training process. Extensive experiments conducted on 11 different datasets demonstrate the effectiveness of our approach, achieving 2.37% and 1.14% average improvement in class-incremental and task-incremental settings, respectively. Code will be available at https://github.com/DeepMed-Lab-ECNU/DPeCLIP.
Figure 1: (a) Comparisons of two approaches for solving the ODCL-CIL task: The first uses all seen categories for classification, while the second selects the corresponding categories for classification through a Task-ID discriminator. (b) Comparisons of Task-ID classification accuracy between MoE-Adapters and ours. (c) Comparisons of our and other methods in ODCL-CIL task. The ODCL task requires evaluating both seen and unseen datasets. The black dashed vertical line indicates the test dataset has not been trained on, and it is assessed through the model’s zero-shot capability.
@@ Introduction
Continual learning (CL) (Dhar et al. 2019; Rebuffi et al. 2017) aims to enable models to continually acquire new knowledge. In recent years, CL has gained significant attention and been applied in various fields, including autonomous driving and medical diagnosis. However, traditional continual learning methods are largely limited to single-domain data, failing to effectively address the complexities of real-world scenarios. In practice, data often come from multiple domains, and there may be significant differences between these domains.
@@
To address this issue, open-domain continual learning (ODCL)  (Zheng et al. 2023; Li et al. 2024) has been explored by many recent researchers (Zheng et al. 2023; Yu et al. 2024; Li et al. 2024). The ODCL task utilizes Vision-Language Model (VLM) (Jia et al. 2021; Li et al. 2022; Radford et al. 2021) to linearly learn data from different domains, acquiring new domain knowledge while preventing the forgetting of old domain knowledge. Due to the zero-shot classification capability of VLM, ODCL is tested on both seen and unseen domains, necessitating the preservation of VLM’s original zero-shot classification abilities. Similar to traditional continual learning, the ODCL task includes both class-incremental (CIL) and task-incremental (TIL) settings. Compared to ODCL-TIL, ODCL-CIL lacks access to the current image’s Task-ID during inference.
@@
ODCL task faces two significant challenges: (1) Two types of Forgetting: Unlike traditional CL task, ODCL task faces not only the forgetting of old knowledge but also the forgetting of zero-shot capabilities. (2) Influence of Category-relatedness Between Domains: For example, although “wheelchair” in Caltech101 and “chair” in CIFAR100 belong to entirely different categories, their semantic similarity may cause confusion in the ODCL-CIL task, leading to a decrease in dataset performance.
@@
Currently, there are two primary technical approaches to addressing ODCL task, as illustrated in Figure 1(a). For the first approach, during the inference stage, the model utilizes all seen categories to classify the test image regardless of which domain it comes from. ZSCL (Zheng et al. 2023) uses a large reference dataset and distills original CLIP knowledge. However, fully fine-tuning CLIP still leads to significant forgetting. ColeCLIP (Li et al. 2024) seeks to minimize interference from similar categories across different domains by maintaining a vocabulary and employing a carefully crafted task prompt. However, it uses the same text embedding for identical categories across datasets, which leads to inconsistencies due to the lack of updates to the old task prompt, despite employing momentum to slow changes in text embeddings. As shown in Figure 1(c), with the progression of training stages, ColeCLIP exhibits significant catastrophic forgetting issues. These methods try to alleviate the forgetting problem, but since the model cannot access data from multiple domains simultaneously during training, and similar categories from different domains may cause confusion during testing, addressing the category-relatedness problem becomes very challenging.
@@
The second approach involves first determining the Task-ID of the test image using a Task-ID discriminator, and then classifying it using the category set corresponding to the identified Task-ID. For example, MoE-Adapters (Yu et al. 2024) proposes a combination of Auto-Encoder (AE) and AlexNet (Krizhevsky, Sutskever, and Hinton 2012) to identify the Task-ID. Subsequently, it explores the performance of CLIP on the ODCL task by adopting a Mixture of Experts (MoE) (Jacobs et al. 1991) structure.Despite the expert freezing mechanism, the continual updating of experts still results in knowledge forgetting. As shown in Figure 1(b) and (c), the effectiveness of the Task-ID discriminator directly affects the model’s final performance.
@@
Compared to the first approach, the second approach provides a better solution to the category-relatedness issue, as a more robust Task-ID discriminator can effectively tackle this problem. However, even with a well-designed task prompt, using cross-domain information does not alleviate the forgetting issue due to inconsistencies between text embeddings and old task prompts. To address category-relatedness and mitigate forgetting of old domains, two factors should be considered: (1) accurately identifying the Task-ID of a test image and using the corresponding category set, and (2) preserving knowledge relevant to each domain. Additionally, the model must retain the original knowledge of CLIP to maintain strong zero-shot capability.
@@
Based on the above considerations, we propose a simple yet effective solution: leveraging intra-domain category-aware prototypes for ODCL in CLIP (Radford et al. 2021) framework, dubbed as Domain Prototype enhanced CLIP (DPeCLIP). To maintain the original zero-shot capability of CLIP, we propose to average the original outputs of CLIP’s image and text modalities belonging to the same category within a domain as prototypes. We propose to (1) distinguish the Task-ID of a test image, and (2) preserve the knowledge only related to each domain, where prototypes are the key to bridging them. Firstly, we propose a training-free Task-ID discriminator method, by utilizing prototypes as classifiers for identifying Task-IDs. Compared to MoE-Adapters, our method significantly improves Task-ID judgment accuracy without introducing additional training parameters. Secondly, to maintain the knowledge corresponding to each domain, we incorporate intra-domain category-aware prototypes as domain prior prompts into training process. For the text branch, we propose a text self-attention module which encodes the relationships of categories within a domain as prompts into the text branch. For the image branch, unlike the text branch, we introduce an image cross-attention module which uses instance-level image embeddings to query category-level prototypes. This ensures that the instance prompt carries information about the relationships between various categories, enabling the learned prompt to encapsulate both domain and instance information.
@@
  • We propose a training-free Task-ID discriminator method that utilizes domain-specific, category-aware prototypes as classifiers to effectively distinguish test images from the original domain.
  • We incorporate intra-domain category-aware prototypes as domain prior prompts into the training process to maintain the knowledge corresponding to each domain.
  • Through comprehensive experiments on 11 datasets, we demonstrate our method achieves state-of-the-art performance in both ODCL-CIL and ODCL-TIL settings, with 4.90% and 3.33% improvement in Last and Forgetting metrics for ODCL-CIL task, compared to the $2$nd best.
@@
To address the issue of catastrophic forgetting, continual learning can be broadly categorized into three primary approaches: replay-based methods (Bang et al. 2021; Chaudhry et al. 2018; Rebuffi et al. 2017; Lopez-Paz and Ranzato 2017; Prabhu, Torr, and Dokania 2020; Shin et al. 2017), regularization-based methods (Kirkpatrick et al. 2017; Zenke, Poole, and Ganguli 2017; Li and Hoiem 2017; Dhar et al. 2019; Douillard et al. 2020; Hou et al. 2019), and parameter expansion-based methods (Gao et al. 2023; Smith et al. 2023; Wang et al. 2022a; Yan, Xie, and He 2021; Zhou et al. 2022; Wang et al. 2022b, c; Wang, Huang, and Hong 2022). Replay-based methods maintain a buffer of previously encountered data and leverage this replay buffer during the learning of new data to retain the original knowledge. Regularization-based methods aim to preserve prior knowledge by constraining the direction of model updates. Parameter expansion-based methods mitigate forgetting by expanding the model’s architecture to accommodate new tasks. Recently, several parameter-efficient fine-tuning (PEFT) based on parameter expansion methods have emerged, such as Low-Rank Adaptation (LoRA) (Hu et al. 2021), prompt-based methods (Lester, Al-Rfou, and Constant 2021), and adapter methods (Houlsby et al. 2019). DualPrompt (Wang et al. 2022b) further mitigates catastrophic forgetting by employing both general and specific prompts. CODA-Prompt (Smith et al. 2023) extends this approach by increasing the scale of prompts and utilizing more prompts simultaneously through a weighting mechanism. LAE (Gao et al. 2023) explores the use of three PEFT methods by implementing both online and offline update strategies.
@@
Currently, numerous studies focus on the performance of CLIP in continual learning, which can be divided into two categories: traditional continual learning (Ding et al. 2022; Zhou et al. 2023; Wang, Huang, and Hong 2022) and ODCL tasks (Zheng et al. 2023; Yu et al. 2024; Li et al. 2024).
Traditional approaches using CLIP (Ding et al. 2022; Zhou et al. 2023; Wang, Huang, and Hong 2022) still focus on a single domain, with the model being simply replaced by CLIP, which does not meet the needs of real life.
ZSCL (Zheng et al. 2023), as the first method proposed for open-domain tasks, balances newly learned knowledge and old knowledge through feature-level distillation with an external dataset. MoE-Adapters (Yu et al. 2024) improve performance in ODCL task by training corresponding AE for Task-ID identification and using a MoE structure. CoLeCLIP (Li et al. 2024) extends open-domain tasks, by maintaining a vocabulary and designing skillfully task prompts to store task-specific knowledge, but this approach inevitably leads to performance degradation in the ODCL-CIL task.
Figure 2: Domain Prototype enhanced CLIP (DPeCLIP) framework, consisting of three stages: (a) Prototype Calculation, where category-aware prototypes are extracted using the original CLIP. (b) Training, where we propose Text Self-Attention (TSA) and Image Cross-Attention (ICA) to provide domain prior prompts, with prototypes as input. (c) Inference, where we use the prototypes to determine Task-IDs for test images and employ the corresponding domain components for classification.
@@
CLIP is a model endowed with zero-shot classification capabilities, comprising an image encoder $\mathcal{I}(\cdot)$ and a text encoder $\mathcal{T}(\cdot)$. To employ the CLIP model for classification tasks, an image $x$ is first processed through the image encoder to extract image features. The target classes $\{c^{1},c^{2},\cdots,c^{J}\}$ for classification are converted into corresponding textual descriptions $\{\hat{c}^{1},\hat{c}^{2},\cdots,\hat{c}^{J}\}$ using a template such as “a photo of a <category>”, where $J$ represents the total number of categories. These textual descriptions are then fed into CLIP text encoder to generate text features. The cosine similarity between the text features and the image features is computed, and the candidate category with the highest similarity is selected as the classification result.
@@
For the ODCL tasks, the model is required to sequentially learn from $N$ distinct datasets (domains). The primary objective of this task is to enhance the recognition of new categories while maintaining the ability to recognize previously learned categories. Given $N$ datasets, $\{\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{N}\}$, as the model learns these datasets linearly, after learning the $n$th dataset $\mathcal{D}_{n}$, it is expected to show good classification capability on the previously learned datasets $\{\mathcal{D}_{1},\mathcal{D}_{2},\cdots,\mathcal{D}_{n}\}$. At the same time, the model aims to preserve its zero-shot classification ability for the unlearned datasets $\{\mathcal{D}_{n+1},\mathcal{D}_{n+2},\cdots,\mathcal{D}_{N}\}$.
To evaluate the model’s performance, we consider two inference settings: task-incremental (ODCL-TIL) and class-incremental (ODCL-CIL). In ODCL-TIL, the Task-ID is known during inference, enabling the use of the corresponding category set for the test image. In contrast, ODCL-CIL does not provide access to the Task-ID, requiring the use of all seen categories for classification.
@@
Figure 2 illustrates the overall architecture of our method. Our method consists of three parts: (a) Prototypes Calculation Stage, (b) Training Stage, and (c) Inference Stage. In the Prototypes Calculation Stage, we utilize the CLIP image encoder to extract image features of training images belonging to the $j$th category within $n$th domain and average them to obtain the mean image feature $\bar{I}_{n}^{j}$. Likewise, we use the CLIP text encoder to extract text features, combined with CLIP templates. By averaging these text features, we obtain the average text feature $\bar{T}_{n}^{j}$. Finally, by performing element-wise summation of $\bar{I}_{n}^{j}$ and $\bar{T}_{n}^{j}$, we derive one intra-domain category-aware prototype $P_{n}^{j}$.
During the training stage, we propose domain prior prompts through two sub-modules: Text Self-Attention (TSA) and Image Cross-Attention (ICA), conditioned on the prototypes as input. In addition, we also learn the shared information for each domain through the general prompts. In this process, both domain prior prompts and the general prompts are updated using cross-entropy loss.
During the inference stage, we first utilize the domain prototypes to determine the Task-ID. Then, based on the identified Task-ID, we select the corresponding TSA, ICA modules, general prompts, and the category set associated with the Task-ID from the domain prior prompt pool, which contain all trained components, to classify the test image.
@@
To obtain intra-domain category-aware prototypes for ODCL, we extract prototypes from the training set using the original CLIP model, as shown in Figure 2(a). Specifically, for the $j$th category in $\mathcal{D}_{n}$, containing $M$ training images, the image prototype $\bar{I}_{n}^{j}$ can be calculated via: $\bar{I}_{n}^{j}=\frac{\sum_{m=1}^{M}\mathcal{I}(x_{n}^{j,m})}{M}$, where $x_{n}^{j,m}$ means the $m$th image belonging to the $j$th category of $\mathcal{D}_{n}$.
We adopt different templates provided by CLIP to generate various textual descriptions for a category name. For the $j$th category in the $n$th domain $c_{n}^{j}$, the textual description generated by the $z$th template is denoted as $\hat{c}_{n}^{j,z}$, the templates come from the original CLIP. Then the text prototype $\bar{T}_{n}^{j}$ is calculated by: $\bar{T}_{n}^{j}=\frac{\sum_{z=1}^{Z}\mathcal{T}(\hat{c}_{n}^{j,z})}{Z}$, where $Z$ represents the total number of templates.
Then, we obtain the intra-domain category-aware prototypes $P_{n}^{j}$ by element-wise summation of $\bar{I}_{n}^{j}$ and $\bar{T}_{n}^{j}$.
@@
Moreover, we propose a category-aware domain prior prompt for the text encoder.Conditioned on the prototype $P_{n}=[{P_{n}^{1}}^{\top},{P_{n}^{2}}^{\top},\cdots,{P_{n}^{S}}^{\top}]^{\top}$ as input, where $P_{n}\in{\mathbb{R}^{S\times d^{\mathcal{T}}}}$. We propose a Text Self-Attention (TSA) module to generate the domain prior prompt $v_{p}$, $v_{p}\in{\mathbb{R}^{S\times d^{\mathcal{T}}}}$. The TSA module is depicted in Figure 2 (with Add and Norm omitted). We use the domain prototype $P_{n}$ to obtain ${Q}$, ${K}$, and ${V}$ in TSA through the formula $Q=W_{Q}\cdot P_{n},K=W_{K}\cdot P_{n},V=W_{V}\cdot P_{n}$, thereby obtaining $v_{p}$.
@@
These domain prior prompts $v_{p}$ replace the general prompt in the $h$-th layer of the text encoder. Thus, the output of the $h$th text encoder layer is:
@@
As shown in Figure 2(b), similar to the text branch, we use the general prompt to learn the general information of the domain for image encoder $\mathcal{I}(\cdot)$. In addition, we utilize the ICA module to generate the instance-level domain prior prompt, thereby providing both domain and instance information simultaneously. Specifically, the image passes through the patch embedding layer to obtain image tokens $[cls^{(1)},i_{1}^{(1)},\cdots,i_{l^{\prime}}^{(1)}]$, where $cls^{(1)}$ is the class token in the first layer and subscript $l^{\prime}$ is the number of image tokens. In the first $h-1$ layers of the image encoder, we concatenate a general prompt $e_{g}$, where $e_{g}\in{\mathbb{R}^{1\times d^{\mathcal{I}}}}$ with the image tokens at each layer, and $d^{\mathcal{I}}$ means the dimension of image token. For example, for the $h-1$ layer $\mathcal{I}^{(h-1)}(\cdot)$ of image encoder, the general prompt $e_{g}^{(h-1)}$ is used by following:
$\begin{split}&[cls^{(h)},t_{1}^{(h)},\cdots,t_{l^{\prime}}^{(h)},e_{g}^{(h)}]=\\ &\mathcal{I}^{(h-1)}([cls^{(h-1)},t_{2}^{(h-1)},\cdots,t_{l^{\prime}}^{(h-1)},e_{g}^{(h-1)}])\vspace{-0.3em}.\end{split}$ (5)